find the x and y intercepts and graph the equation plotting the intercepts please show all of your work. 6x-7y=42
Answer by mananth(16946)

equation 6x-7y=42
7y=6x-42
y=(6/7)x+6

when x= 0 y= 6 ( 0 , 6 )
when y = 0 x= -7 ( -7 , 0 )
y intercept =6
X intercept = -7